A hollow cylindrical tube open at both ends is made of iron 3 cm thick. if the external diameter be 56 cm and the length of the tube be 182 cm,find the number of cubic cmin it?

Grade:8

1 Answers

Aabid Hussain
7 years ago
Hi Vineet,
Thank you for asking
Height = 182 cm
External Radius = 56/2 = 28 cm
Internal Radius = 56-6/2 = 50/2 = 25 cm
Total material used = πh(R2 - r2)
= 22/7 × 182 × 159
= 90948 cm3